OpenStudy (morgan_home.study):

Which property would allow you to use mental computation to simplify the problem 27 + 15 + 3 + 5?

OpenStudy (welshfella):

commutative property bring the 27 and 3, and 15 and 5 together;- 27 + 3 + 15 + 5 = 30 + 20 = 50

OpenStudy (mrnood):

 so the ORDER of addition is not important: What is that property called?